In what two positions will a converging lens of focal length 7.5 cm form images of an object on a screen 40 cm from the object?


1
Expert's answer
2021-02-28T07:28:05-0500
"\frac{1}{f}=\frac{1}{v}+\frac{1}{u}\\\frac{1}{7.5}=\frac{1}{40-v}+\frac{1}{v}\\v_1=10\ cm\\v_2=30\ cm"